# FinCon: A Synthesized LLM Multi-Agent System with Conceptual Verbal Reinforcement for Enhanced Financial Decision Making

### v2

@@ 5.2.1 Single Asset Trading Task
5.2.2 Portfolio Management Task
To address the first research question, we compare FinCon’s performance with the Markowitz Mean-Variance (MV) portfolio [78] and FinRL [76] in managing a small portfolio of TSLA, MSFT, and PFE. Those assets are chosen by the stock selection agent from 42 stocks with sufficient amount of news data (over $800$ new messages in the training and test periods together), referring to Figure 9 in Appendix 7.2. The training & testing periods and backbone algorithm are consistent with those used in the single-asset trading task. For the Markowitz MV portfolio, we estimate the covariance matrix and expected returns using the same training data. For FinRL, we use 5-year training data before the test period. Our results, detailed in Table 3, show that FinCon outperforms both the Markowitz MV portfolio and FinRL, achieving significantly higher cumulative returns and Sharpe Ratio, and lower Maximum Drawdown (MDD).
However, there are stronger indications of hallucination in the multi-asset decision-making task compared to single-asset trading due to increased complexity and input length. Although FinCon reduces the workload of individual agents by distributing tasks and focusing on essential investment insights, it can still occasionally produce incorrect information, such as generating non-existent memory IDs and misclassifying action types (e.g., outputting ‘buy’ and ‘sell’ instead of ‘long’ and ‘short’). Handling multi-asset decision-making tasks requires intricate logic and substantial market information, posing a significant challenge for LLMs to comprehend extended contexts. This challenge has left portfolio management an untapped field for previous language agent studies. However, FinCon demonstrates the potential of constructing agent systems that can address complex financial problems through resource optimization, even when managing a relatively compact portfolio.
[…]
Table 3: Key performance metrics comparison among all portfolio management strategies. FinCon leads all performance metrics.
Figure 5: Portfolio value (CR for portfolio) changes over time for all the strategies.
5.3 Ablation Studies
In this section, we comprehensively evaluate our unique risk control component through two ablation studies. The first study assesses its ability to control risk within episodes using conditional value at risk (CVaR). The second study highlights the importance of the over-episode risk control mechanism in updating the trading manager agent’s beliefs for a holistic understanding of current trading circumstances. Consistency in training and testing periods is maintained with the main experiments for both single stock trading and portfolio management tasks.
5.3.1 The Effectiveness of Within-Episode Risk Control mechanism via CVaR
@@
Our results demonstrate that implementing CVaR in FinCon is highly effective across all financial metrics for both task types, as shown in Table 4 and Fig 6. For single-asset trading tasks, FinCon without within-episode risk control yields negative CRs and significantly higher MDDs, underperforming compared to the Buy-and-Hold strategy (CR of GOOG: $22.42\%$, CR of NIO: $-51.057\%$), highlighting the severe consequences of ignoring environmental risks. In portfolio management, the CR increases dramatically from $16.994\%$ to $121.018\%$ with within-episode risk control, demonstrating its effectiveness in risk supervision even amid non-uniform market trends.
[…]
@@
Figure 6: CRs of FinCon with vs. without implementing CVaR for within-episode risk control show that the CVaR mechanism significantly improves FinCon’s performance. This is evident from two metrics: (a) cumulative returns over time for single stocks in both bullish and bearish market conditions, and (b) portfolio value over time for a multi-asset portfolio. In both cases, FinCon with CVaR demonstrates substantially higher cumulative returns.
@@
5.3.2 The Efficacy of Over-Episode Belief Updates Using CVRF
@@
The results in Table 5 and Figure 7 indicate that the over-episode belief update mechanism is more critical than within-episode risk control in enhancing FinCon’s decision-making. Without this functionality, key metrics like CR, SR, and MDD are lower than without the within-episode risk control in single asset trading. Although the CR of $20.677\%$ slightly outperforms the Equal-Weighted ETF strategy’s $19.983\%$ for portfolio management, the SR and MDD are both lower. However, with the belief update feature, performance significantly improves, achieving a CR of $121.018\%$ and an SR of $3.435$. These results demonstrate that using CVRF to update investment beliefs over episodes efficiently steers the agent’s investment beliefs towards more profitable directions. FinCon achieves superior performance on multiple tasks, with learning gains evident after just four training episodes, requiring far fewer episodes than traditional RL algorithmic trading agents.
[…]
Table 5: Key metrics FinCon with vs. without implementing belief updates for over-trajectory risk control. The performance of FinCon with the implementation of CVaR won a leading performance in both single-asset trading and portfolio management tasks.
@@
Figure 7: CRs of FinCon with vs. without belief updates for over-trajectory risk control. (a) The CRs over time for single stocks. The performance of FinCon with belief updates consistently leads in both bullish and bearish market conditions. (b) The CRs over time for multi-asset portfolio. FinCon’s performance with belief updates also won a substantially higher CR.
@@
6 Conclusion
In this paper, we present FinCon, a novel LLM-based multi-agent framework for financial decision-making tasks, including single stock trading and portfolio management. Central to FinCon is the Synthesized Manager-Analyst hierarchical communication structure and a dual-level risk control component. This communication method channels financial data from multiple sources to specialized analyst agents, who distill it into key investment insights. The manager agent then synthesizes these insights for decision-making. Our experimental evaluations demonstrate the efficacy of our risk control mechanism in mitigating investment risks and enhancing trading performance. Additionally, the streamlined communication structure reduces overhead. The dual-layered risk control component introduces a novel approach to defining agent personas, enabling dynamic updates of systematic risk and market beliefs within agent communication. A valuable future research direction would be to scale FinCon’s framework to manage large-sized portfolios comprising tens of assets, while maintaining the impressive decision-making quality demonstrated with smaller portfolios. Given the LLM’s input length constraint, a critical challenge lies in striking an optimal balance between information conciseness through agent distillation and potential performance deterioration when extending the current context window. Addressing this will be essential for ensuring quality-assured outcomes.

### v3

@@ A.6 Single Stock Trading Result Graphs
A.7 Detailed Ablation Study
A.7.1 The Effectiveness of Within-Episode Risk Control mechanism via CVaR
@@
Our results demonstrate that implementing CVaR in FinCon is highly effective across all financial metrics for both task types, as shown in Table 4 and Fig 6. For single-asset trading tasks, FinCon without within-episode risk control yields negative CRs and significantly higher MDDs, underperforming compared to the Buy-and-Hold strategy (CR of GOOG: $22.42\%$, CR of NIO: $-77.210\%$), highlighting the severe consequences of ignoring environmental risks. In portfolio management, the CR increases dramatically from $14.699\%$ to $113.836\%$ with within-episode risk control, demonstrating its effectiveness in risk supervision even amid non-uniform market trends.
@@
Figure 6: CRs of FinCon with vs. without implementing CVaR for within-episode risk control show that the CVaR mechanism significantly improves FinCon’s performance. This is evident from two metrics: (a) cumulative returns over time for single stocks in both bullish and bearish market conditions, and (b) portfolio value over time for a multi-asset portfolio. In both cases, FinCon with CVaR demonstrates substantially higher gains.
@@
A.7.2 The Efficacy of Over-Episode Belief Updates Using CVRF
@@
The results in Table 5 and Figure 7 indicate that the over-episode belief update mechanism is more critical than within-episode risk control in enhancing FinCon’s decision-making. Without this functionality, key metrics like CR, SR, and MDD are lower than without the within-episode risk control in single asset trading. Although the CR of $28.432\%$ outperforms the Equal-Weighted ETF strategy’s $9.344\%$ for portfolio management, the SR of $1.181$ is higher than Equal-Weighted ETF strategy’s $0.492$, with the belief update feature, performance significantly further improves. It can achieve a CR of $113.836\%$ and an SR of $3.269$. These results demonstrate that using CVRF to update investment beliefs over episodes efficiently steers the agent’s investment beliefs towards more profitable directions. FinCon achieves superior performance on multiple tasks, with learning gains evident after just four training episodes, requiring far fewer episodes than traditional RL algorithmic trading agents.
@@
Figure 7: CRs of FinCon with vs. without belief updates for over-episode risk control. (a) The CRs over time for single stocks. The performance of FinCon with belief updates consistently leads in both bullish and bearish market conditions. (b) The portfolio values over time for multi-asset portfolio. FinCon’s performance with belief updates also won a substantially higher gains.
@@
A.8 Raw Data Sources
We assessed the performance of FinCon using multi-modal financial data from January 3, 2022, to June 10, 2022, sourced from reputable databases and APIs including Yahoo Finance (via yfinance), Alpaca News API, and Capital IQ, detailed explained in Table. These data, initially stored in the Raw Financial Data Warehouse as available observations of the financial market environment, are diverged into the corresponding FinCon’s Analysts’ Procedural Memory Databases based on timeliness through working memory’s summarization operation.
Data Sources News data associated with ticker: News data is sourced from REFINITIV REAL-TIME NEWS mainly contains news from Reuters. Form 10-Q, Part 1 Item 2 (Management’s Discussion and Analysis of Financial Condition and Results of Operations): Quarterly reports (Form 10-Q) are required by the U.S. Securities and Exchange Commission (SEC). Form 10-k, Section 7 (Management’s Discussion and Analysis of Financial Condition and Results of Operations): Annual reports (Form 10-K) are required by the U.S. Securities and Exchange Commission (SEC), sourced from EDGAR, and downloaded via SEC API. Historical stock price: Daily open price, high price, close price, adjusted close price, and volume data from Yahoo Finance. Zacks Equity Research: Zacks Rank: The Zacks Rank is a short-term rating system that is most effective over the one- to three-month holding horizon. The underlying driver for the quantitatively determined Zacks Rank is the same as the Zacks Recommendation and reflects trends in earnings estimate revisions. Zacks Analyst: Reason to Sell, Reason to Buy, and potential risks. Earning Conference Calls (ECC): ECC is a type of unstructured financial data (audio) that is crucial for understanding market dynamics and investor sentiment. The company executive board delivers ECC about recent financial outcomes, future projections, and strategic directions. Recent studies have underscored the importance of not only the textual content of these calls but also the audio feature. Analyses have revealed that the audio elements—such as tone, pace, and inflections—offer significant predictive value regarding company performance and stock movements [94, 95, 96].
Table 6: Raw data and memory warehouses of FinCon
A.9 Distribution of Data
Figure 9: The distribution of news from REFINITIV REAL-TIME NEWS for the 42 stocks in the experiments
Figure 10: The distribution of 10k10q from Securities and Exchange Commission (SEC) for the 42 stocks in the experiments
Figure 11: The distribution of Analyst Report from Zacks Equity Research for the 42 stocks in the experiments
A.10 Formulas of Classic Financial Metrics for Risk Estimator and Decision-making Task Performance Evaluation
The risk estimator uses the following metrics:
Profit and Loss (PnL)[97]: PnL quantifies the net outcome of trading activities over a specified period by accounting for the realized gains and losses from financial instruments like stocks and derivatives.
Value at Risk (VaR) of PnL[97]: VaR is a statistical tool used to estimate the potential loss in a portfolio, within a defined confidence interval. Mathematically, it is defined as Equation 3:
$\text{VaR}_{\alpha}(PnL)=\inf\left\{l\in\mathbb{R}:\mathbb{P}(PnL\leq l)\geq\alpha\right\}$ (3)
where $\alpha$ is the confidence level.
Conditional Value at Risk (CVaR) of PnL[97]: CVaR is a statistical tool used to estimate the expected potential loss worse than the VaR value in a portfolio, within a defined confidence interval. Mathematically, it is defined as Equation 4:
$\text{CVaR}_{\alpha}(PnL)=\mathbb{E}\Big{\{}PnL|PnL\leq\text{VaR}_{\alpha}(PnL)\Big{\}}$ (4)
where $\alpha$ is the confidence level.
The performance evaluation of algorithmic trading agents incorporates the following metrics:
Cumulative Return of PnL [90]: Cumulative Return is a key trading performance metric because it provides a comprehensive insight into investment performance, especially for strategies that emphasize long-term growth and reinvestment. The effectiveness of different investment strategies is evaluated based on their Cumulative Returns, which reflect the total change in value over time. In this study, we compute Cumulative Returns over the specified period by summing daily logarithmic returns, as outlined in Equation 5. This method is widely accepted in the finance area due to its ability to precisely capture minor price fluctuations and symmetrically address gains and losses. In essence, a higher Cumulative Return typically indicates a more effective strategy.
Cumulative Return $=\sum_{t=1}^{n}r_{i}$
$=\sum_{t=1}^{n}\left[\ln\left(\frac{p_{t+1}}{p_{t}}\right)\cdot\text{action}_{t}\right],$ (5)
where $r_{i}$ represents the PnL for day $t+1$, $p_{t}$ is the closing price on day $t$, $p_{t+1}$ is the closing price on day $t+1$, and $\text{action}_{t}$ denotes the trading decision made by the model for that day.
Portfolio Value: Portfolio value represents the total worth of all the investments held in a portfolio at a given point in time. It is a metric used only in the portfolio management task.
$\textbf{Cumulative Simple Return}_{t}=\prod_{k=1}^{t}(1+\textbf{Daily Simple Return}_{t})-1$ (6)
$\textbf{Portfolio Value}_{t}=\textbf{Initial Investment Amount}\times(1+\textbf{Cumulative Simple Return}_{t})$ (7)
, where the initial amount is set as \$$\$1,000,000$.
Sharpe Ratio of PnL [91]: Sharpe Ratio is another core metric for evaluating investment performance and adjusting returns for risk. It is calculated by dividing the portfolio’s average PnL ($R_{p}$) over the risk-free rate ($R_{f}$) by its volatility ($\sigma_{p}$), as shown in Equation 8. This metric adjusts returns for risk, with a higher ratio indicating better risk-adjusted performance. Essential in comparing different portfolios or strategies, it contextualizes performance against similar investments. Although a Sharpe Ratio above 1 is typically considered favorable and above 2 as excellent, these benchmarks can vary depending on the context of comparison.
Sharpe $\textbf{Sharpe Ratio}=\frac{R_{p}-R_{f}}{\sigma_{p}}$ (8)
Max Drawdown of PnL [92]: Max Drawdown is a metric for assessing risk. It represents the most significant decrease in a portfolio’s value, from its highest ($P_{\text{peak}}$) to its lowest point ($P_{\text{trough}}$) until a new peak emerges, detailed in Equation 9. Indicative of investment strategy robustness, a smaller Max Drawdown suggests reduced risk.
Max $\textbf{Max Drawdown}=\text{max}(\frac{P_{\text{peak}}-P_{\text{trough}}}{P_{\text{peak}}})$ (9)
A.11 Baseline and Comparative Models on Single Stock Trading Task
Buy-and-Hold strategy (B&H):
A passive investment approach, where an investor purchases stocks and holds onto them for an extended period regardless of market fluctuations, is commonly used as a baseline for comparison of stock trading strategies.
LLM trading agents:
We evaluate FinCon against four LLM agents in the context of stock trading.
  • General-purpose Generative Agents – GA:The generative AI agent by Park et al. [20], originally intended to simulate realistic human behavior and make everyday decisions, has been adapted here for specific stock trading tasks. This agent’s architecture includes a memory module that employs recency, relevance, and importance metrics to extract pivotal memory events for informed decision-making. However, it does not provide a layered memory module to effectively differentiate the time sensitivities unique to various types of financial data. Additionally, although it features a profiling module to define agent attributes like professional background, the model does not specify the agent’s persona. In our experiments, we modified the original prompt template created by Park et al., which was intended for general daily tasks, to suit financial investment tasks.
  • FinGPT: A novel open-source LLM framework specialized for converting incoming textual and numeric information into informed financial decision-making, introduced by Yang et al[31]. It claims superiority over the traditional buy-and-hold strategy.
  • FinMem:FinMem employs a specialized profiling module and self-adaptive risk settings for enhanced market robustness. Its memory module integrates working memory and layered long-term memory, enabling effective data processing. This allows FinMem to leverage market insights and improve trading decisions [32].
  • FinAgent:FinAgent developed upon FinMem, which leverages the use of tool-using capabilities of LLMs to incorporate multi-modal financial data [33]. It claims an further improved trading performance on single asset trading (stocks and cryptocurrencies).
DRL trading agents:
As the FinMem is practiced and examined on the basis of single stock trading and discrete trading actions, we choose three advanced DRL algorithms fitting into the same scenarios according to the previous and shown expressive performance in the work of Liu et al [98, 99]. The DRL training agents only take numeric features as inputs.
  • Advantage Actor-Critic (A2C): A2C ([100]) is applied to optimize trading actions in the financial environment. It operates by simultaneously updating both the policy (actor) and the value (critic) functions, providing a balance between exploration and exploitation.
  • Proximal Policy Optimization (PPO): PPO ([101]) is employed in stock trading due to its stability and efficiency. One salient advantage of PPO is that it maintains a balance between exploration and exploitation by bounding the policy update, preventing drastic policy changes.
  • Deep Q-Network (DQN): DQN ([102]) is an adaptation of Q-learning, that can be used to optimize investment strategies. Unlike traditional Q-learning that relies on a tabular approach for storing Q-values, DQN generalizes Q-value estimation across states using deep learning, making it more scalable for complex trading environments.
A.12 Portfolio Management
Markowitz Portfolio Selection [1]: introduced by Harry Markowitz in 1952, is a framework for constructing portfolios that optimize expected return for a given level of risk or minimize risk for a given level of expected return. This method uses expected returns, variances, and covariances of asset returns to determine the optimal asset allocation, thereby balancing risk and return through diversification.
FinRL-A2C [48]: is an RL algorithm proposed to address single stock trading and portfolio optimization problems in Liu et al.. The RL models make trading decisions (i.e., portfolio weights) based on the observation of previous market conditions and the brokerage information of the RL agents. The implementation of this algorithm  [^2] https://github.com/AI4Finance-Foundation/FinRL-Meta is provided and is used as baselines in our study.
Equal-Weighted ETF [103]: is a portfolio giving equal allocation to all stocks, similar to a buy-and-hold strategy in single-stock trading, can provide a benchmark on market trends.
A.13 Ranking Metrics for Procedural Memory in FinCon
Upon receiving an investment inquiry, each agent in FinCon retrieves the top-$K$ pivotal memory events from its procedural memory, where $K$ is a hyperparameter. These events are selected based on their information retrieval score. For any given memory event $E$, its information retrieval score $\gamma^{E}$ is defined by
$\gamma^{E}=S_{\text{Relevancy}}^{E}+S_{\text{Importance}}^{E}$ (10)
which is adpated from Park et al [20] but with modified relevancy and importance computations, and is scaled to $[0,1]$ before summing up. Upon the arrival of a trade inquiry $P$ in processing memory event $E$ via LLM prompts, the agent computes the relevancy score $S_{\text{Relevancy}}^{E}$ that measures the cosine similarity between the embedding vectors of the memory event textual content $\mathbf{m_{E}}$ and the LLM prompt query $\mathbf{m_{P}}$, which is defined as follows:
$S_{\text{Relevancy}}^{E}=\frac{\mathbf{m_{E}}\cdot\mathbf{m_{P}}}{\|\mathbf{m_{E}}\|_{2}\times\|\mathbf{m_{P}}\|_{2}}$ (11)
Note that the LLM prompt query inputs trading inquiry and trader characteristics. On the other hand, the importance score $S_{\text{Importance}}^{E}$ is inversely correlates with the time gap between the inquiry and the event’s memory timestamp $\delta t=t_{\text{P}}-t_{E}$, mirroring Ebbinghaus’s forgetting curve [104]. More precisely, if we denote the initial score value of memory event $v^{E}$ and degrading ratio $\theta\in(0,1)$, then the importance score is computed via
$S_{\text{Importance}}^{E}=v^{E}\times\theta^{\delta t}$ (12)
Note that the ratio $\theta$ measures the diminishing importance of an event over time, which is inspired by design of [20]. But in our design, the factors of recency and importance are handled by one equation. Different agents in FinCon admit different choices of $\{v^{E},\theta\}$ for memory event $E$.
Additionally, an access counter function facilitates memory event augmentation, so that critical events impacting trading decisions can be augmented by FinCon, while trivial events are gradually faded. This is achieved by using the LLM validation tool Guardrails AI [105] to track critical memory ID. A memory ID deemed critical to investment gains receives $+5$ to its importance score $S_{\text{Importance}}^{E}$. This access counter implementation enables FinCon to capture and prioritize crucial events based on type and retrieval frequency.
A.14 Detailed Configurations in Experiments
The training period was chosen to account for the seasonal nature of corporate financial reporting and the duration of data retention in FinCon’s memory module. The selected training duration ensures the inclusion of at least one publication cycle of either Form 10-Q, ECC, or Form 10-K. This strategy ensures that the learned conceptualized investment guidance considers a more comprehensive scope of factors. Additionally, the training duration allowed FinCon sufficient time to establish inferential links between financial news, market indicators, and stock market trends, thereby accumulating substantial experience. Furthermore, we set the number of top memory events retrieved for each agent at 5. We ran FinCon. The reported performance outcomes are based on the setting that achieved the highest cumulative return during the testing phase.
To maintain consistency in the comparison, the training and testing phases for the other three LLM-based agents were aligned with those of FinMem. For parameters of other LLM-based agents that are not encompassed by FinMem’s configuration, they were kept in accordance with their original settings as specified in their respective source codes.
FinCon’s performance was benchmarked against that of the most effective comparative model, using Cumulative Return and Sharpe Ratio as the primary evaluation metrics. The statistical significance of FinCon’s superior performance was ascertained through the non-parametric Wilcoxon signed-rank test, which is particularly apt for the non-Gaussian distributed data.
A.15 FinCon performance on extreme market conditions
To further illustrate the robustness of FinCon’s performance, we assess its effectiveness in two distinct scenarios: (1) a single-asset trading task using TSLA and (2) a portfolio management task involving a combination of TSLA, MSFT, and PFE. Our evaluation focuses on key financial metrics, including Cumulative Returns (CRs), Sharpe Ratios (SRs), and Maximum Drawdown (MDD). The training period spanned from January 17, 2022, to March 31, 2022, while the testing phase covered April 1, 2022, to October 15, 2022. This specific timeframe was chosen due to the elevated levels of the CBOE Volatility Index (VIX), which averaged above 20, signaling greater market volatility during these months.
As demonstrated in Table 7 and Figure LABEL:subfig:single_asset_cr, FINCON is the sole agent system to achieve positive Cumulative Returns (CRs) and Sharpe Ratios (SRs) in single stock trading tasks. Regarding portfolio management tasks, the results of all baselines (four benchmarks) are detailed in Table 8 and Figure LABEL:subfig:portfolio_management_cr. In these comparisons, FINCON consistently attained the highest scores in the primary performance metrics.
[…]
Table 7: Key performance comparison for single asset trading under the high volatility condition using TSLA as an example. FinCon leads all performance metrics.
Figure 12: CR changes over time across all the strategies under the high volatility condition using TSLA as an example of the single asset trading task.
[…]
Table 8: Key performance comparison among all portfolio management strategies of Portfolio1 under the high volatility condition. FinCon leads all performance metrics.
Figure 13: Portfolio1 value changes over time for all the strategies under the high volatility condition.